# Cost-Free Personalization via Information-Geometric Projection in Bayesian Federated Learning

### v1

@@ 4 Personalization via Information-Geometric Projection
The main contribution of this work lies in the interpretation of the personalization problem as a projection problem. Given a divergence function, the objective is to project the global posterior $p_{g}$ onto the projection set of the $k^{th}$ client, defined as the sphere centered in the local posterior $p_{k}$ and of radius $r_{k}$. Conceptually, the projection can be understood as identifying the distribution within the “local neighborhood of the client’s posterior that best aligns with the global model. In this setting, the radius $r_{k}$ encodes the degree of personalization required by the client, where smaller values of $r_{k}$ imply a stricter alignment to the local posterior rather than to the global model, as shown in Figure 1. In the following, we define a local sphere $\mathcal{S}_{k}$ under a divergence function $D$.
@@
(Local Sphere $\mathcal{S}_{D}(p,r)$) Given a statistical manifold $\mathcal{M}$ and a divergence function $D$, the local sphere $\mathcal{S}_{D}(p,r)$ for $p\in\mathcal{M}$ and $r\in[0,\infty)$a as the set
$\mathcal{S}_{D}(p,r)=\{q\in\mathcal{M}:\quad D(q||p)\leq r\}\subseteq\mathcal{M}$
To simplify the notation, we will indicate with $\mathcal{S}_{k}=\mathcal{S}_{D}(p_{k},r_{k})$ the local sphere of the $k^{th}$ client centered in the local posterior $p_{k}$ and of radius $r_{k}$. We are now ready to provide a formal definition of the personalization problem as a projection problem.
@@
(Projection on $\mathcal{S}_{k}$) Given a statistical manifold $\mathcal{M}$, a divergence function $D$, a global posterior $p_{g}\in\mathcal{M}$, and a local sphere $\mathcal{S}_{k}$, we define the projection problem of $p_{g}$ on $\mathcal{S}_{k}$ as
@@
We identify the solution to this projection problem as the personalized posterior distribution $p_{g,k}=\arg\min_{p\in\mathcal{S}_{k}}D(p||p_{g})$ of client $k$.
@@
By varying the radius $r_{k}\in[0,D(p_{k}||p_{g})]$, the solutions of the associated projection problem define the geodesic between $p_{k}$ and $p_{g}$, i.e., the shortest possible path connecting the two distributions.
To support our derivation, we begin by presenting the definition of a barycenter with respect to a given divergence $D$.
@@
The following mild assumption is crucial for establishing the equivalence between the projection and barycenter problems, as shown in Theorem 1.
@@
The divergence metric is a convex function in the first argument.
@@
Most commonly used divergences, including the family of $f$-divergences and Wasserstein-$p$ distances, are convex in both of their arguments.
@@
Given Assumption 2, the solution of the projection problem (8) is equivalent to the weighted barycenter problem (2), i.e.,
$p_{g,k}=p^{*}_{D}(\{p_{g},p_{k}\},\{w_{g},w_{k}\})$ (10)
where $w_{g}$ and $w_{k}$ are defined as
@@
There exists an inverse proportionality between the Lagrangian multiplier $\lambda$ and the radius $r_{k}$. Specifically, as $\lambda\to 0$, the radius $r_{k}\to\infty$, corresponding to the case where the personalized posterior aligns with the global posterior. Conversely, in the limit $\lambda\to\infty$, the radius $r_{k}$ vanishes ($r_{k}\to 0$), leading to the assignment of the local posterior to the personalized posterior. Therefore, by selecting a $\lambda$ we implicitly fix the personalization radius $r_{k}$.
The main advantage of the equivalence between projections and barycenters lies in the greater tractability of the barycentic formulation. For example, under Assumption 1, analytical solutions exist for both the RKL divergence and the Wasserstein-2 distance, as shown in [31, 2]. These closed-form solutions enable a straightforward and computationally cost-free personalization process.
Figure 2: Effect of $\lambda$ on the performance of local versus global data distributions. Results are reported for the CIFAR-10 dataset. Notably, $\lambda=0$ corresponds to the global model and $\lambda\to\infty$ corresponds to the local model.
@@ 5.1 Experimental Setting
Datasets and Heterogeneity Simulation
We evaluate our approach on three widely used image classification benchmarks: FashionMNIST, SVHN, and CIFAR-10. To simulate the label shift across the 10 clients, we consider Dirichlet-based partitioning, where, similarly to prior works [33, 50, 47, 48, 35, 39], we sample client-specific label distributions from a Dirichlet distribution with concentration parameter $\beta=0.5$, inducing non-i.i.d data across clients.
Model Architecture and Hyperparameters
@@
Our model architecture consists of two convolutional layers with 5×5 kernels and ReLU activations, each followed by a 2×2 max-pooling layer. The extracted features are flattened and passed through three fully connected layers of sizes 120, 84, and 10, respectively, to produce the final class logits. For all state-of-the-art methods we compare against, we closely adhere to the implementation details provided in the original papers. We note that the complexity of our architecture is similar to the models used in the selected state-of-the-art methods, ensuring a fair and consistent comparison. Whenever possible, we use the official codebases for implementation; otherwise, we carefully reproduce the setups following the authors’ guidelines to maintain alignment with their reported settings. We make an exception for pFedBayes on FashionMNIST, where, instead of the original Multi-Layer Perceptron model with one hidden layer, we use our architecture to produce a fair comparison.
Personalization Step
@@
For personalization, we restrict our experimental evaluation to the RKLB and WB, as both barycenters preserve the Gaussian structure of the distributions. This preservation is a critical design choice, as it facilitates the adaptability of the personalized models. In particular, it ensures that the posterior distribution of the personalized model remains aligned with those of both the global and local models, which are assumed to be Gaussian under the variational learning framework.
@@
Table 1: Comparison of all methods across four evaluation settings: personalized models on local data (PM on LD), personalized models on global data (PM on GD), global models on local data (GM on LD), and global models on global data (GM on GD). Results are reported for three datasets (FashionMNIST, SVHN, CIFAR-10) using Accuracy, NLL, and ECE as evaluation metrics.
@@ 5.2 Comparison between the different aggregation methods considered
Our experiments investigate different combinations of global update methods (EAA, RKLB, WB) and personalization techniques (RKLB, WB). Statistical tests, provided in Appendix C, support our observation of comparable performance across these configurations. To reduce redundancy and enhance readability, we limit the results presented in the subsequent experiments to WB for both global updates and personalization.
We analyze the impact of the personalization parameter $\lambda$ on model performance across three datasets under heterogeneity simulated with the Dirichlet distribution. Across all datasets, we observe a consistent trade-off between performance on global and local data. Results on CIFAR-10 are reported in Figure 2. The global data, i.e., the union of all client test sets, is approximately uniform over classes while the local data follows different Dirichlet distributions. Notably, $\lambda=0$ coincides with the global model, and $\lambda\to\infty$ corresponds to the local model. As $\lambda$ increases, performance on the global distribution deteriorates in terms of accuracy, calibration (ECE), and uncertainty quantification (measured by NLL), whereas local performance improves and remains stable over a certain range of $\lambda$.. The increased ECE and NLL for local models ($\lambda\to\infty$) on global data suggest that over-personalization may reduce generalization and model confidence, particularly for underrepresented classes. Conversely, at lower values of $\lambda$, the model achieves improved global performance but fails to effectively capture client-specific distributions and shows reduced confidence on local data. These results underscore the importance of $\lambda$ in controlling the trade-off between generalization and personalization, enabling the model to adapt effectively to heterogeneous data distributions in non-i.i.d. federated settings.
5.4 Comparison with state-of-the-art
In this section, we provide a detailed analysis of the results presented in Table 1 and Figure 3. All reported values correspond to the mean and standard deviation over three independent runs.
@@ 5.4 Comparison with state-of-the-art
Figure 3: Trade-offs between local and global performance for personalized models. Each subplot shows results for a different evaluation metric: Accuracy (left), ECE (center), and NLL (right). Points represent method–dataset pairs. For accuracy, the top-right indicates a better performance trade-off; for ECE and NLL, the bottom-left is preferable. Our method (with $\lambda=1$) consistently achieves a favorable balance across all metrics and datasets.
Results of Personalized Models
@@
It is worth noting that all methods in the comparison, except FedAVG, employ personalized models. Consequently, we compare the local models of FedAVG, i.e., the models sent by the clients before aggregation, to the personalized models produced by the other methods, including ours. The results are presented in the first two groups of rows in Table 1 and Figure 3.
@@
  • On Local Data: Personalized models are designed to adapt to client-specific distributions, which is clearly reflected in their local accuracy scores, shown in the first group of rows in Table 1. Our method achieves performance comparable to or better than the baselines; for example, it achieves the highest accuracy in SVHN and is second in CIFAR-10. In terms of calibration (ECE) and uncertainty quantification (NLL), our method consistently yields low error rates on local data, achieving the best performance in SVHN and second best in CIFAR-10.
@@
  • On Global Data: A key limitation of many personalized methods is the degradation in global performance due to overfitting to client-specific distributions. This trade-off is particularly evident in FedPop on FashionMNIST, where it achieves the best local accuracy, but performs poorly on global evaluations. In contrast, our method maintains strong generalization, achieving the highest global accuracy across all datasets: 84.61% on FashionMNIST, 79.87% on SVHN, and 57.75% on CIFAR-10. These represent improvements of 3.87%, 11.20%, and 5.70%, respectively, over the second-best method on each dataset. Furthermore, our method consistently achieves the lowest ECE and NLL on global data, demonstrating both well-calibrated predictions and effective uncertainty quantification.
@@
  • Trade-off Analysis: Figure 3 visualizes the trade-off between local and global performance for personalized models in the three metrics. For accuracy, the top-right region indicates favorable performance; for ECE and NLL, better performance lies in the bottom left. Our method consistently appears closest to the optimal region in all three plots, demonstrating strong local accuracy without sacrificing global generalization. The plots confirm that our method achieves a superior balance across all datasets, combining high accuracy with well-calibrated and confident predictions.
Results of Global Models
@@
We compare the performance of global models across all methods, excluding FedPop, which by design does not maintain a global model. The results discussed below correspond to the last two groups of rows in Table 1. The results for global models on both local and global data are closely similar, as the GM-on-LD setting reports the average performance of the global model evaluated across all clients’ local data. Our method performs strongly in the global setting, achieving the best results in SVHN and CIFAR-10, across all three metrics: accuracy, ECE, and NLL. This demonstrates not only high predictive performance but also well-calibrated and confident uncertainty estimates, highlighting the robustness of our approach even in non-personalized settings.

### v2

@@ 4 Personalization via Information-Geometric Projection
The main contribution of this work lies in interpreting the personalization problem as an information-geometric projection problem. Given a divergence function $D$, the objective is to project the global posterior $p_{g}$ onto the projection set of the $k^{th}$ client, defined as a sphere centered at the local posterior $p_{k}$ with radius $r_{k}$. The projection is performed after FL training has been completed. Conceptually, this projection can be viewed as identifying the distribution within the “local neighborhood” of the client’s posterior that best aligns with the global model. The radius $r_{k}$ quantifies the degree of personalization required by client $k$, where smaller values of $r_{k}$ correspond to stronger adherence to the local posterior, while larger values allow greater influence from the global model, as shown in Figure 1. In the following, we formally define the local sphere $\mathcal{S}_{k}$ induced by the divergence function $D$.
@@
(Local Sphere $\mathcal{S}_{D}(p,r)$) Given a statistical manifold $\mathcal{M}$ and a divergence function $D$, the local sphere $\mathcal{S}_{D}(p,r)$ centered at $p\in\mathcal{M}$ with radius $r\in[0,\infty)$ is defined as the set
$\mathcal{S}_{D}(p,r)=\{q\in\mathcal{M}:\quad D(q||p)\leq r\}\subseteq\mathcal{M}.$
To simplify the notation, we denote by $\mathcal{S}_{k}=\mathcal{S}_{D}(p_{k},r_{k})$ the local sphere associated with the $k^{th}$ client, centered at the local posterior $p_{k}$ with radius $r_{k}$. We are now ready to formally define the personalization problem as a projection problem.
@@
(Projection onto $\mathcal{S}_{k}$) Given a statistical manifold $\mathcal{M}$, a divergence function $D$, a global posterior $p_{g}\in\mathcal{M}$, and a local sphere $\mathcal{S}_{k}$, we define the projection of $p_{g}$ onto $\mathcal{S}_{k}$ as the optimization problem
@@
The solution to this projection problem is identified as the personalized posterior distribution of client $k$, denoted by $p_{g,k}=\arg\min_{p\in\mathcal{S}_{k}}D(p||p_{g})$.
@@
By varying the radius $r_{k}\in[0,D(p_{g}||p_{k})]$, the solutions of the associated projection problem trace the geodesic between $p_{k}$ and $p_{g}$, i.e., the shortest possible path connecting the two distributions under the geometry induced by the divergence $D$.
To support our derivation, we first introduce the definition of a barycenter with respect to a divergence $D$.
@@
The following mild assumption is essential for establishing the equivalence between the projection and barycenter formulations, as shown in Theorem 1.
@@
The divergence metric is a convex function in its first argument.
@@
Most divergences used in practice, such as the family of $f$-divergences and the Wasserstein-$p$ distances, are convex in both arguments, and therefore satisfy Assumption 2.
@@
Under Assumption 2, the solution of the projection problem 1 is equivalent to that of the weighted barycenter problem 2, i.e.,
$p_{g,k}=p^{*}_{D}(\{p_{g},p_{k}\},\{w_{g},w_{k}\}),$ (9)
where the weights $w_{g}$ and $w_{k}$ are given by
@@
There exists an inverse proportional relationship between the Lagrangian multiplier $\lambda$ and the radius $r_{k}$. Specifically, as $\lambda\to 0$, the radius $r_{k}\to\infty$, corresponding to the case where the personalized posterior coincides with the global posterior. Conversely, in the limit $\lambda\to\infty$, the radius $r_{k}$ vanishes ($r_{k}\to 0$), implying that the personalized posterior collapses to the local posterior. Hence, by selecting a value of $\lambda$, we implicitly determine the personalization radius $r_{k}$.
The main advantage of the equivalence between projections and barycenters lies in the improved tractability of the barycentic formulation. For instance, under Assumption 1, analytical solutions are available for both the RKL divergence and the Wasserstein-2 distance, as shown in Koliander et al. (2022); Álvarez-Esteban et al. (2016). These closed-form expressions enable a straightforward and computationally efficient personalization procedure, incurring virtually no additional cost.
@@ 5.1 Experimental Setting
Datasets and Heterogeneity Simulation.
Interpretability and reliable uncertainty quantification are crucial in high-stakes domains such as banking and healthcare. When these institutions participate in FL, they typically do so in a cross-silo setting, where the number of clients is small but each holds a large local dataset. Reflecting this application focus, we concentrate our empirical evaluation on cross-silo simulations rather than cross-device benchmarks Caldas et al. (2018). To simulate these conditions, we evaluate our approach on three widely used image classification benchmarks: FashionMNIST, SVHN, and CIFAR-10, and we induce label shift across 10 clients using Dirichlet-based partitioning. Following prior works Li et al. (2020); Yurochkin et al. (2019); Wang et al. (2020a; b); Lin et al. (2020); Ozer et al. (2022), we sample client-specific label distributions from a Dirichlet distribution with concentration parameter $\beta=0.5$, which induces label shift and yields non-i.i.d. data across clients.
Metrics.
Following the taxonomy of evaluation metrics in Table IV of Tan et al. (2023), our evaluation focuses mainly on model performance and trustworthy AI criteria. For model performance, we report overall global-model accuracy as well as the average and variance of personalized-model accuracies, together with the worst-performing 10% of clients’ accuracy. The latter two also connect to the fairness column in Table IV, and we elaborate on this in Section 5.5. In addition, our Bayesian formulation explicitly targets uncertainty quantification and model calibration, which are not explicitly listed among the model-performance or trustworthy-AI metrics in Tan et al. (2023). We therefore report Expected Calibration Error (ECE) and Negative Log-Likelihood (NLL) alongside accuracy-based measures to assess both the quality and the reliability of the predictive distributions. In contrast, we do not report system-performance metrics such as fault tolerance or system scalability, since our method has similar per-round complexity to FedAvg (see Section 3), and our focus here is on the statistical behavior of the personalized posteriors rather than on system-level aspects.
Models Architecture and Hyperparameters.
@@
Our model architecture consists of two convolutional layers with 5×5 kernels and ReLU activations, each followed by a 2×2 max-pooling layer. The extracted features are flattened and passed through three fully connected layers of sizes 120, 84, and 10, respectively, to produce the final class logits. For all state-of-the-art methods we compare against, we closely follow the implementation details provided in the original papers. The complexity of our architecture is comparable to that of the models used in these methods, ensuring a fair and consistent comparison. Whenever possible, we use the official codebases for implementation; otherwise, we carefully reproduce the setups following the authors’ guidelines to maintain alignment with their reported settings. We make an exception for pFedBayes on FashionMNIST, where, instead of the original multi-layer perceptron with one hidden layer, we use our convolutional architecture to enable a fair comparison.
It is worth noting that training with IVON can be sensitive to hyperparameter settings, particularly to the initialization of the Hessian. Therefore, we report in Table 1 the hyperparameters used for training on each dataset. Here, $N_{k}$ denotes the size of the training dataset for the $k^{th}$ client.
Table 1: IVON Hyperparameters.
params FashionMNIST SVHN CIFAR-10
initial learning rate 0.1 0.1 0.1
final learning rate 0.01 0.01 0.01
weight decay 2e-4 2e-4 2e-4
batch size 64 64 64
ESS $N_{k}$ $N_{k}$ $N_{k}$
initial hessian ($h_{0}$) 5 2 1
MC sample while training 1 1 1
MC sample while testing 10 10 10
Personalization Step.
@@
For personalization, we restrict our experimental evaluation to RKLB and WB, as both barycenters preserve the Gaussian structure of the distributions. This preservation is a critical design choice, as it facilitates the adaptability of the personalized models. In particular, it ensures that the posterior distribution of the personalized model remains consistent with those of both the global and local models, which are assumed to be Gaussian under the variational learning framework.
Our experiments investigate various combinations of global update methods (EAA, RKLB, WB) and personalization techniques (RKLB, WB). Statistical tests, presented in Appendix B, support our observation that these configurations yield comparable performance. To reduce redundancy and enhance readability, we therefore restrict the results reported in the subsequent experiments to the WB method for both global updates and personalization.
We analyze the impact of the personalization parameter $\lambda$ on model performance across three datasets under heterogeneity simulated using the Dirichlet distribution. Across all datasets, we observe a consistent trade-off between performance on global and local data. Results on CIFAR-10 are reported in Figure 2 and the results on FashionMNIST and SVHN are reported in Appendix C. The global data, i.e., the union of all client test sets, are approximately uniform across classes while the local data follow distinct Dirichlet distributions. Notably, $\lambda=0$ corresponds to the global model, whereas $\lambda\to\infty$ represents the local model. As $\lambda$ increases, performance on the global distribution deteriorates in terms of accuracy, calibration (ECE), and uncertainty quantification (measured by NLL), whereas local performance improves and remains stable within a certain range of $\lambda$. The increased ECE and NLL for local models ($\lambda\to\infty$) on global data suggest that over-personalization may reduce generalization and model confidence, particularly for underrepresented classes. Conversely, at lower values of $\lambda$, the model achieves improved global performance but fails to effectively capture client-specific distributions, showing reduced confidence on local data. These results underscore the importance of $\lambda$ in controlling the trade-off between generalization and personalization, allowing the model to adapt effectively to heterogeneous data distributions in non-i.i.d. federated settings.
Figure 2: Effect of $\lambda$ on performance across local and global data distributions. Results are reported for the CIFAR-10 dataset. Notably, $\lambda=0$ corresponds to the global model, whereas $\lambda\to\infty$ corresponds to the local model.
5.4 Comparison with State-of-the-Art
In this section, we provide a detailed analysis of the results presented in Table 2 and Figure 3. All reported values correspond to the mean and standard deviation computed over three independent runs.
@@ 5.4 Comparison with State-of-the-Art
Table 2: Comparison of all methods across four evaluation settings: personalized models on local data (PM on LD), personalized models on global data (PM on GD), global models on local data (GM on LD), and global models on global data (GM on GD). Results are reported for three datasets (FashionMNIST, SVHN, and CIFAR-10) using Accuracy, NLL, and ECE as evaluation metrics.
@@
Figure 3: Trade-offs between local and global performance for personalized models. Each subplot presents results for a different evaluation metric: Accuracy (left), ECE (center), and NLL (right). Points represent method–dataset pairs. For Accuracy, the top-right region indicates a better performance trade-off, whereas for ECE and NLL, the bottom-left region is preferable. Our method (with $\lambda=1$) consistently achieves a favorable balance across all metrics and datasets.
Results of Personalized Models.
@@
It is worth noting that all methods in the comparison, except FedAvg, employ personalized models. Consequently, we compare the local models of FedAvg, i.e., the models sent by the clients before aggregation, with the personalized models produced by the other methods, including ours. The results are presented in the first two groups of rows in Table 2 and Figure 3.
@@
  • On Local Data: Personalized models are designed to adapt to client-specific distributions, which is clearly reflected in their local accuracy scores, shown in the first group of rows in Table 2. Our method achieves performance comparable to or better than the baselines; for example, it attains the highest accuracy on SVHN and the second-highest on CIFAR-10. In terms of calibration (ECE) and uncertainty quantification (NLL), our method consistently yields low error rates on local data, achieving the best performance on SVHN and second-best on CIFAR-10.
@@
  • On Global Data: A key limitation of many personalized methods is the degradation in global performance due to overfitting to client-specific distributions. This trade-off is particularly evident in FedPop on FashionMNIST, which achieves the best local accuracy but performs poorly on global evaluations. In contrast, our method maintains strong generalization, achieving the highest global accuracy across all datasets: 84.61% on FashionMNIST, 79.87% on SVHN, and 57.75% on CIFAR-10. These represent improvements of 3.87%, 11.20%, and 5.70%, respectively, over the second-best method on each dataset. Furthermore, our method consistently achieves the lowest ECE and NLL on global data, demonstrating both well-calibrated predictions and effective uncertainty quantification.
@@
  • Trade-off Analysis: Figure 3 visualizes the trade-off between local and global performance for personalized models across the three metrics. For accuracy, the top-right region indicates favorable performance; for ECE and NLL, better performance lies in the bottom-left region. Our method consistently appears closest to the optimal region in all three plots, demonstrating strong local accuracy without sacrificing global generalization. The plots confirm that our method achieves a superior balance across all datasets, combining high accuracy with well-calibrated and confident predictions.
Results of Global Models.
@@
We compare the performance of global models across all methods, excluding FedPop, which by design does not maintain a global model. The results discussed below correspond to the last two groups of rows in Table 2. The results for global models on both local and global data are closely aligned, as the GM-on-LD setting reports the average performance of the global model evaluated across all clients’ local data. Our method performs strongly in the global setting, achieving the best results on SVHN and CIFAR-10 across all three metrics: accuracy, ECE, and NLL. This demonstrates not only high predictive performance but also well-calibrated and confident uncertainty estimates, highlighting the robustness of our approach even in non-personalized settings.
5.5 Client-Level Fairness Analysis
Figure 4: Performance of the global model on clients’ local data. Top row: average accuracy. Bottom row: worst-case performance. Each column corresponds to a different dataset.
Despite the growing interest in BFL, its fairness implications remain a largely overlooked aspect. Although not the main focus of this work, we take a first step toward a systematic investigation of client-level fairness from an accuracy-parity perspective. In Appendix D.1, we provide an extended literature review that motivates our choice of fairness metrics.
As a comprehensive fairness evaluation requires comparing shared models across clients under consistent conditions, we focus on the performance of the global model. It is important to note that FedPop does not produce a global model. Although it achieves strong performance in terms of personalized models, the absence of a single shared model renders it unsuitable for evaluating fairness from a global perspective. As a result, the notion of fairness across all clients is not applicable in its context.
Our results, presented in Figure 4, show that the global model produced by our method is comparable to state-of-the-art approaches in terms of both mean accuracy and worst 10% accuracy across clients. Additionally, we conduct a study to examine the effect of Bayesian layers on algorithmic fairness in federated learning, and report our findings in Appendix D.2.
6 Broader Applicability of the Proposed Method
Although the use of information geometry in BFL is natural, our method also applies to any optimization problem defined over probability spaces rather than point estimates. The approach can naturally extend to domain adaptation scenarios: when a model is trained on source data $A$ and new target data $B$ become available, retraining only on $B$ may lead to catastrophic forgetting of the source domain. Our projection framework provides a principled trade-off, retaining prior knowledge while integrating new information. This flexibility is further enhanced by the use of IVON, which combines the efficiency of deterministic learning with the ability to perform inference and adaptation in the posterior space. However, it is important to note that our approach fundamentally differs from standard domain adaptation methods based on optimal transport Courty et al. (2016), which operate on data distributions rather than model posteriors. Appendix E provides additional insights and discussion.